The median listing price of a home in Hollywood in 2010 was $348 thousand. In 2018 the
median listing price is $625 thousand. Using the information given answer the following questions since
2010.
a. Find the linear equation of the median listing price of a home in Hollywood since 2010. State the
equation in slope-intercept form and use exact values.
b. Using the linear equation found in (a) if the trend continues what will the median listing price of a
home in Hollywood be in 2030? State your solution in a complete sentence and round your solution to
the first decimal place.
c. In what year will the median listing price of a home in Hollywood be $902 thousand? State your
solution in a complete sentence and round to the nearest year.

Answer by Boreal(15235)

let 2010 be x=0
so we have (0, 348)
and we have (8, 625)
the slope is 277/8 or 34.625
the point slope formula is y-y1=m(x-x1), m slope (x1, y1) point
y-348=34.625*x
y==34.625x+348
in 2030, x=20
y=34.625(20)+348
=$1,040,500
902=34.625x+348
554=34.625x
x=16 years after 2010
2026 is the answer.